# Phase dynamics and their role determining energy flux in hydrodynamic shell models

### v1

@@ V Conclusions
The surprisingly simple description of the triad phase dynamics, and its direct prediction of the alignment and energy flux, begs the question of its extension to more realistic contexts and flows. Although we have seen that the conditions of scale-independent triad phase statistics approximately hold in the inertial ranges of full 3D-like shell model simulations, this is not the case in 2D-like full model simulations, as we have already noted in the main text. This is due to the restriction in the inertial ranges of full systems requiring a constant-in-scale energy flux. That said, our simplified analytical model of the triad phase dynamics may still hold without having to invoke a power law energy spectrum and $n$-independent triad phase dynamics – predictions of the triad phase behavior could depend on $n$. Also, an investigation of how the model’s noise properties change between full and phase-only shell model runs is worth considering. Preliminary results from the two full shell model runs mentioned in this work reveal hints that the noise has a larger variance and larger tails in its PDF, but that otherwise the same properties hold.
@@
A model of triad phase dynamics and their implication on the flux of conserved quantities for realistic flows could improve our understanding of established results in turbulence, but could also be helpful in understanding the behavior that deviates from what is expected. In the context of anisotropic turbulent flows, such as in geophysical and astrophysical systems, the energy can cascade both up- and down-scale in what is known as a bidirectional energy cascade [4]. The bidirectional cascade exhibits striking behavior as the relevant geophysical parameter (which produces anisotropy) is varied: starting from a fully forward cascade, the bidirectional cascade seems to appear via a bifurcation at a precise parameter value. Neither this bifurcation nor the functional dependence of the inverse-to-forward cascade ratio with the control parameter are well understood. We expect that similar triad phase dynamics models can shed light on the dynamical and physical mechanisms responsible for this behavior.
@@ Acknowledgements.
The authors would like to thank Alexandros Alexakis, Luca Biferale, Michele Buzzicotti, Javier Jiménez and Adrian van Kan for helpful discussions during the development of this project and manuscript. S.J.B. has received funding from the European Union’s Horizon 2022 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101109237. The code used to run the shell model simulations described in the main text is v1.0Z of ‘MHD shell model’ [82]. It is preserved at https://doi.org/10.5281/zenodo.15800546, available via the Creative Commons Attribution 4.0 International (CC-BY) license. The Figshare repository https://doi.org/10.6084/m9.figshare.29469659 contains (i) full data sets for shell model runs; (ii) temporal data for the two runs used in Figure 1; (iii) the scripts used to generate the figures in the main text.
@@ Appendix A Colored harmonic noise
where $\tau\coloneq\Gamma/\Omega_{0}^{2}$. The steady-state PDF for these dynamics can be found using the Fokker-Planck equation, resulting in,
$PDF(\theta)=\frac{1}{\mathcal{Z}}\left[1+\tau\mathcal{K}\sin(\theta)-\frac{\mathcal{K}(\Gamma\sin(\theta)+\mathcal{K})}{(\Gamma+\mathcal{K}\sin(\theta))}\right]\exp{\left[\frac{\mathcal{K}}{D}\sin(\theta)-\frac{\tau\mathcal{K}^{2}}{2D}\cos(\theta)^{2}-\frac{\Gamma^{2}-\mathcal{K}^{2}}{D(\Gamma+\mathcal{K}\sin(\theta))}-\frac{\Gamma}{D}\log(\Gamma+\mathcal{K}\sin(\theta))\right]}.$ (25)

### v2

@@ V Conclusions
The surprisingly simple description of the triad phase dynamics, and its direct prediction of the alignment and energy flux, begs the question of its extension to more realistic contexts and flows. Although we have seen that the conditions of scale-independent triad phase statistics approximately hold in the inertial ranges of full 3D-like shell model simulations, this is not the case in 2D-like full model simulations, as we have already noted in the main text and in Appendix B. This is due to the restriction in the inertial ranges of full systems requiring a constant-in-scale energy flux. That said, our simplified analytical model of the triad phase dynamics may still hold without having to invoke a power law energy spectrum and $n$-independent triad phase dynamics – predictions of the triad phase behavior could depend on $n$. Also, an investigation of how the model’s noise properties change between full and phase-only shell model runs is worth considering. Preliminary results from the two full shell model runs mentioned in this work reveal hints that the noise has a larger variance and larger tails in its PDF, but that otherwise the same properties hold.
@@
A model of triad phase dynamics and their implication on the flux of conserved quantities for realistic flows could improve our understanding of established results in turbulence, but could also be helpful in understanding the behavior that deviates from what is expected. In the context of anisotropic turbulent flows, such as in geophysical and astrophysical systems, the energy can cascade both up- and down-scale in what is known as a bidirectional energy cascade [24, 74, 1, 67, 8, 57]. The bidirectional cascade exhibits striking behavior as the relevant geophysical parameter (which produces anisotropy) is varied: starting from a fully forward cascade, the bidirectional cascade seems to appear via a bifurcation at a precise parameter value [72, 71, 11, 65, 1, 78, 79]. Neither this bifurcation nor the functional dependence of the inverse-to-forward cascade ratio with the control parameter are well understood. We expect that similar triad phase dynamics models can shed light on the dynamical and physical mechanisms responsible for this behavior.
@@ Acknowledgements.
The authors would like to thank Alexandros Alexakis, Luca Biferale, Michele Buzzicotti, Javier Jiménez and Adrian van Kan for helpful discussions during the development of this project and manuscript. S.J.B. has received funding from the European Union’s Horizon 2022 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101109237. The code used to run the shell model simulations described in the main text is v1.0Z of ‘MHD shell model’ [10]. It is preserved at https://doi.org/10.5281/zenodo.15800546, available via the Creative Commons Attribution 4.0 International (CC-BY) license. The Figshare repository [9], preserved at https://doi.org/10.6084/m9.figshare.29469659, contains (i) full data sets for shell model runs; (ii) temporal data for the two runs used in Figure 1; (iii) the scripts used to generate the figures in the main text.
@@ Appendix A Colored harmonic noise
where $\tau\coloneq\Gamma/\Omega_{0}^{2}$. The steady-state PDF for these dynamics can be found using the Fokker-Planck equation, resulting in
$PDF(\theta)$ $=$ $\frac{1}{\mathcal{Z}}\left[1+\tau\mathcal{K}\sin(\theta)-\frac{\mathcal{K}(\Gamma\sin(\theta)+\mathcal{K})}{(\Gamma+\mathcal{K}\sin(\theta))}\right]$ (25)
$\times\exp{\left[\frac{\mathcal{K}}{D}\sin(\theta)-\frac{\tau\mathcal{K}^{2}}{2D}\cos(\theta)^{2}-\frac{\Gamma^{2}-\mathcal{K}^{2}}{D(\Gamma+\mathcal{K}\sin(\theta))}-\frac{\Gamma}{D}\log(\Gamma+\mathcal{K}\sin(\theta))\right]}.$
@@
Appendix B Comparison to 2D-like ‘full’ shell model runs
In this Appendix, we briefly discuss 2D-like shell model runs of the non-phase-only (i.e., ‘full’) shell model, Eq. (1). We performed 11 runs varying $\gamma$ from $\gamma=0.13$ to $\gamma=3.4$. All runs have $N=35$ modes, $\lambda=1.5$, $k_{0}=\lambda^{-15}$, and $a=1$. Both small-scale and large-scale dissipation were included, so that $D_{n}=-\nu k_{n}^{2}-\mu(\delta_{n,0}+\delta_{n,1})$, where $\nu=2.15\times 10^{-4}$ is the viscosity and $\mu=2.10\times 10^{-5}/\gamma$ is the large-scale drag acting only on the first two modes. We scale $\mu$ with $1/\gamma$ arbitrarily, since we found that the large-scale drag needed to prevent condensation at the last two modes decreased with increasing $\gamma$. We used constant-energy-injection-rate forcing, $f_{n}=\varepsilon(\delta_{n,n_{f}}+\delta_{n,n_{f}+1})/(2u_{n}^{*})$, with a forcing mode number $n_{f}=15$ and energy injection rate $\varepsilon=1$. Runs were initialized with random phases and power-law energy spectra with exponent 2/3-2/3. Equation (1) was integrated in time using a fourth order Runge-Kutta method with time step $\Delta t=5\times 10^{-4}$. Statistics were collected once steady state was reached.
Figure 6: Full 2D-like shell model run for $\gamma=1.26$. In (a) the energy spectrum (averaged at steady state) demonstrates the presence of an inertial range with a power-law energy spectrum containing an exponent close to $\alpha=-1.31$, which we fit with a dashed blue line. In (b) we see that the energy flux is purely inverse and the flux of the second conserved quantity is forward. Panel (c) shows the steady-state PDFs of the triad phases at each mode number, and panel (d) shows the corresponding (negative) alignment associated with the PDFs, revealing a strong scale-dependence.
Fig. 6 shows results from the case of $\gamma=1.26$, which is representative of all runs for which $0.7<\gamma<2$. We find an inverse cascading inertial range for $n<n_{f}$, with a constant negative flux of energy on average and an energy spectrum with an approximate power-law exponent of $\alpha=-1.31$ (best fit outside of the dissipation ranges and forcing scales). In this inertial range, like that of real 2D fluid systems, there is no flux of enstrophy. The enstrophy has a forward flux to larger $n$ for $n>n_{f}$, as expected. As seen from the bottom row in Fig. 6, the triad phases are aligned such that $\langle\sin(\theta_{n})\rangle<0$ for all $n$. However, the alignment strength varies by two orders of magnitude over the range of $n$ in the simulation (Fig. 6(d)). In the range $n<n_{f}$, the alignment strength seems to increase as a power law of the wavenumber (exponentially in mode number), although understanding why that is is beyond the scope of this work. As noted in the main text, the combination of $n$-dependence of the alignment strength and the power-law energy spectrum results in a constant energy flux, cf. Fig. 6(b), given by Eq. (II.1).
Between runs, the primary difference observed is the time-averaged kinetic energy power-law exponents for the inverse-cascading inertial range, $\alpha$. Remarkably, the spectral exponents all fall within the regions where inverse energy flux is predicted in our model under the phase-only assumptions (Fig. 7). This is rather surprising, given the fact that our zero-flux lines are based on the assumption that the statistics are $n$-independent, yet all of the full runs have clearly $n$-dependent alignment.
Figure 7: Steady-state energy spectrum exponents of the inverse-cascading inertial range for ‘full’ 2D-like shell model simulations at various values of $\gamma$ plotted over our model’s prediction of the rescaled energy flux for a phase-only run with a given value of $(\alpha,\gamma)$. Despite the $n$-independent phase statistics of the phase-only model, which breaks down in the full shell model simulations, the spectrum exponents all lie within the inverse-cascading regions of the parameter space.
Runs with $\gamma>2$ seem to have somewhat strange behavior in the inertial range for $n<n_{f}$. It contains two distinct regions: one region of modes (at small $n$, or large scales) seemingly at thermal equilibrium with a flat KE spectrum, and a region with non-zero triad phase alignment and a power-law energy spectrum with the measured slopes found in Fig. 7. It seems possible that this is a result of triad phases being influenced by those at $n>n_{f}$, in the enstrophy cascade, which appears to be operating as expected in these runs.
Finally, runs with $\gamma<0.7$ look very similar to those for $0.7<\gamma<2$, except for the peculiar fact that there is a non-zero, sustained inverse enstrophy cascade for $n<n_{f}$.